OpenStudy (pagen13):

Graph the function. Describe the transformation of the parent function f(x)=x^2. f(x)=(x-9)^2+2

OpenStudy (pagen13):

@campbell_st

OpenStudy (pagen13):

@jchick

OpenStudy (misty1212):

HI!!

OpenStudy (misty1212):

do you know what \(y=x^2\) looks like?

OpenStudy (pagen13):

y=4?

OpenStudy (pagen13):

@misty1212

OpenStudy (misty1212):

you lost me the question was, what does the graph of \(y=x^2\) look like

OpenStudy (misty1212):

i think you might be confused as to what the question is asking compared to the graph of \(y=x^2\) the graph of \(y=(x-9)^2+2\)is shifted Left 9 Up 2

OpenStudy (pagen13):

So what does that mean? I am so confused

OpenStudy (jchick):

Sorry can't help right now I will check in a little while

OpenStudy (pagen13):

@jchick ok

OpenStudy (jchick):

Sorry did you get your answer?